OpenStudy (anonymous):

Two resistors connected in series have an equivalent resistance of 767.5 Ω. When they are connected in parallel, their equivalent resistance is 127 Ω. Find the resistance of each resistor.

OpenStudy (tkhunny):

So \(R_{1} + R_{2} = 767.5\) and \(\dfrac{1}{R_{1}}+\dfrac{1}{R_{2}} = \dfrac{1}{127}\)? One should point out that \(R_{1}\;and\;R_{2}\) are interchangeable in these two formulas. Don't be surprised if you get both answers when trying to solve for only one. We call this "symmetry".